OpenStudy (anonymous):

You invested $9000 between two accounts paying 6% and 8% annual interest, respectively. If the total interest earned for the year was $649, how much was invested at each rate? $____ was invested at 6% and $____ was invested at 8%.

OpenStudy (tkhunny):

Name the investments! $S was invested at 6% and $E was invested at 8% Now, translate the pr4oblem statement in terms of S and E. "You invested $9000 between two accounts" "the total interest earned for the year was $649" Go! Show ANY work. Do it!

OpenStudy (anonymous):

(9000-x) (0.06) + x (0.08) = 649 540 - 0.06 x + 0.08 x = 649 0.02 x = 649 - 540 = 109 x = 5450 and (9000-5200) = 3550 checks

OpenStudy (tkhunny):

Of course, actually WRITING DOWN the definition of 'x' would have been an excellent first step.